OpenStudy (anonymous):

6x^3+30x^2+6x+30 completely factor the polynomial

ganeshie8 (ganeshie8):

6x^3+30x^2+6x+30 = 6(x^3 + 5x^2+x+5) =6(x^2(x+5) + x+5) =6(x+5)(x^2+1)

ganeshie8 (ganeshie8):

trick here is to group the terms, and factor out common terms

OpenStudy (anonymous):

got it thanks ganeshie8

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

:)